A car accelerate uniformly at the rate of 4m/5, from an initial velocity of 36km/h from r 30s. find the distance covered during this period


1
Expert's answer
2020-11-12T09:32:33-0500
"s=vt+0.5at^2\\s=\frac{36}{3.6}(30)+0.5(4.5)(30)^2=2325\ m"